Five years ago, Diane secured a bank loan of $380,000 to help finance the purchase of a loft in the San Francisco Bay area. TI just need help with the last two answers. I am not sure why they were counted as incorrect? Thanks.
Answer #1

c)
Rate = 6%/12
Nper = 30*12 = 360
PV = -364344.52
FV = 0
Monthly payment can be calculated by using the following excel formula:
=PMT(rate,nper,pv,fv)
=PMT(6%/12,30*12,-364344.52,0)
= $2184.43

Monthly payment = $2184.43

d)
Reduction in monthly payment = $3057.57 - $2184.43 = $873.14